# VGLD: Visually-Guided Linguistic Disambiguation for Monocular Depth Scale Recovery

### v2

@@ Abstract
We propose a robust method for monocular depth scale recovery. Monocular depth estimation can be divided into two main directions: (1) relative depth estimation, which provides normalized or inverse depth without scale information, and (2) metric depth estimation, which involves recovering depth with absolute scale. To obtain absolute scale information for practical downstream tasks, utilizing textual information to recover the scale of a relative depth map is a highly promising approach. However, since a single image can have multiple descriptions from different perspectives or with varying styles, it has been shown that different textual descriptions can significantly affect the scale recovery process. To address this issue, our method, VGLD, stabilizes the influence of textual information by incorporating high-level semantic information from the corresponding image alongside the textual description. This approach resolves textual ambiguities and robustly outputs a set of linear transformation parameters (scalars) that can be globally applied to the relative depth map, ultimately generating depth predictions with metric-scale accuracy. We validate our method across several popular relative depth models(MiDas, DepthAnything), using both indoor scenes (NYUv2) and outdoor scenes (KITTI). Our results demonstrate that VGLD functions as a universal alignment module when trained on multiple datasets, achieving strong performance even in zero-shot scenarios. Code is available at: https://github.com/pakinwu/VGLD
@@ 1 Introduction
Monocular depth estimation is a fundamental and long-standing task in computer vision, with widespread applications in autonomous driving[41], augmented reality[12], and 3D reconstruction[33]. The goal is to predict a dense depth map from a single RGB image. However, performing 3D reconstruction from a single image is an ill-posed problem because perspective projection leads to a loss of one dimension during image formation, and any point along the projection ray may correspond to the same image coordinate. As a result, the distance from the camera to the scene cannot be directly measured from a single view. In the absence of camera calibration parameters, additional sensor measurements (e.g., IMU[49], LiDAR[29]), or strong priors (e.g., pre-trained depth models), a scale ambiguity problem arises. While multiple images (e.g., stereo images or video) can resolve this by localizing points in space, modern large-scale training models typically use data from diverse sources, with significant differences in data types and distributions. These sources include single RGB images, video streams, RGB images without camera calibration, RGB images with sensor parameters, and so on. The differences in these data settings make solving the scale ambiguity problem particularly crucial.
To address the scale ambiguity, one approach is to train on datasets from multiple domains (e.g., indoor and outdoor)[39, 6, 53, 54], learning depth based on the distribution characteristics of the training set. However, biases introduced by these specific datasets can affect the model’s generalization ability, limiting its transferability across domains[36]. Another approach is to find complementary information or constraints shared by these diverse datasets and use this additional information to alleviate the scale ambiguity problem to some extent. Recent works have begun to use language as a supplemental information source, as language has the practical advantage of not requiring additional expensive synchronized and calibrated sensors (e.g., LiDAR, time-of-flight cameras) for data collection. By leveraging existing publicly available pre-trained image description models and vision-language models[38, 20], we can generate textual descriptions for images from different sources and distributions. RSA[59] is a pioneering work in this innovative approach. RSA[59] posits the hypothesis that language (e.g., text descriptions or captions) can be used to infer the scale of 3D scenes, exploring whether linguistic modalities can resolve scale ambiguity in 3D reconstruction from a single image, and successfully converts scale-less relative depth into metrically-scaled depth based on a general relative depth estimation model[39, 6, 53, 54].
However, purely linguistic descriptions suffer from ambiguity. Experiments show that the same image can have multiple descriptions that are semantically similar or even identical. As shown in Figure 1, the RSA[59] estimator produces different estimates for the same image depending on the different textual input. This linguistic ambiguity significantly impacts RSA’s scale recovery. Despite this, language, as an independent input modality distinct from images, is not influenced by factors such as lighting, object orientation, occlusion, or scene layout, making it a highly promising general-purpose tool.
To stabilize the contribution of the linguistic modality, we incorporate high-level semantic information from the corresponding image, extracted via the CLIP Image Encoder[38], alongside the textual description. This approach applies a Visually-Guided mechanism to the linguistic input, effectively mitigating the issue of Linguistic Disambiguation. Our approach, VGLD, also addresses the problem of large differences in depth distributions across different domains (e.g., indoor vs. outdoor). We draw inspiration from ZoeDepth[5] and propose a Domain Router Mechanism (DRM), which automatically determines the domain of the data and routes it to the appropriate estimation head. This mechanism enables automatic routing to the appropriate scope based on domain input and aligns relative depth with metric scale to predict consistent metric depth across domains, achieving cross-domain generalization. The scale recovery from relative depth maps to absolute depth maps has multiple recovery paths or ways. To optimize the training trajectory of scalars (scale, shift) parameters, we introduce pseudo-labels $(k_{lm},b_{lm})$ obtained from the Levenberg-Marquardt algorithm, a nonlinear optimization method that combines gradient descent with the Gauss-Newton method to solve the least squares problem. Through additional supervised training, we effectively guide the model to follow the correct training trajectory, leading to robust depth scale recovery.
Our contributions are as follows: (i) We integrate high-level semantic information from the corresponding image alongside the textual description, thereby stabilizing the output of the scalars parameters; (ii) We introduce the Domain Router Mechanism, which aids in solving the cross-domain estimation problem; (iii) We leverage the Levenberg-Marquardt algorithm to optimize the training trajectory and guide the model’s training process; (iv) We conduct extensive experiments to evaluate the performance of our method in both indoor and outdoor scenarios, analyzing its sensitivity to text captions and its zero-shot generalization capability.
@@
2 Related Work
2.1 Monocular Depth Estimation
@@
Monocular Depth Estimation (MDE) is a fundamental task in computer vision, with its development generally following two main directions: relative depth estimation[39, 40, 60, 6, 53, 54, 21] and metric depth estimation[9, 10, 50, 3, 4, 30, 28, 1, 7, 25, 31, 37, 40, 47, 52, 56, 17, 29]. The goal of metric depth estimation is to predict pixel-wise depth values in metric units (e.g., meters), and models are typically trained by minimizing the discrepancy between predicted and ground-truth depth maps. In contrast, relative depth estimation focuses on inferring the ordinal relationships between pixel pairs, without providing any information about scale or units. A notable early milestone in this field was Eigen et al.[9], the first to apply Convolutional Neural Networks (CNNs) to MDE. More recent methods such as AdaBins [3], LocalBins[4] and Binsformer[28] reformulate the depth regression problem as a classification task through depth discretization. Multi-task learning strategies have also been explored: GeoNet[37] integrates surface normal estimation, while AiT[34] incorporates instance segmentation, both to enhance depth prediction through joint training. MiDaS[39, 6] and Diversedepth[55] advances relative depth estimation by pretraining on a diverse mixture of datasets, achieving strong generalization across domains. More recently, Depth Anything[53, 54] unleashes the potential of large-scale unlabeled images and multi-dataset training, substantially improving generalization and robustness across a wide range of scenarios. In addition, diffusion-based[19, 11, 21, 48, 62, 45] methods, such as DDP [19], Marigold [21], and GeoWizard [11], adapt powerful diffusion priors to the depth estimation task via fine-tuning, enabling significant performance gains. Despite these advances, metric depth estimation remains highly domain-sensitive[56], often exhibiting poor robustness when applied to unseen environments. Meanwhile, relative depth estimation, being inherently scale-invariant, cannot be directly used for applications requiring metric 3D reconstruction. VGLD serves as a universal alignment framework that converts relative depth into consistent metric estimation, enabling robust performance across diverse domains, including both indoor and outdoor scenes.
2.2 Metric Depth Scale Recovery
Relative depth estimation models have emerged as strong backbones for many metric depth Scale Recovery tasks, owing to their impressive cross-domain generalization and robustness. Building on MiDaS[39], DPT[40] replaces the original convolutional architecture with a Vision Transformer and adapts the model to metric depth through fine-tuning on datasets with absolute scale annotations. ZoeDepth[5] adopts MiDaS[39] as the encoder and introduces a powerful decoder equipped with a metric bins module, demonstrating its ability to transfer to metric depth via fine-tuning on metric datasets. Depth Anything[53] further extends ZoeDepth[5] by replacing the MiDaS[39] encoder with its own model, implicitly transforming relative depth into metric depth. Other approaches, such as Metric3D[17, 56], zeroDepth[16] and UniDepth[36], recover metric scale by leveraging or predicting camera intrinsics. PromptDA[29] introduces a low-cost LiDAR sensor as a prompt to guide the Depth Anything[53] model toward accurate metric depth predictions. Despite their progress, these scale recovery methods are typically restricted to specific domains and often require access to ground-truth depth for scale and shift alignment at test time. To address this, RSA[59] introduces a novel approach that explicitly aligns relative depth to metric scale using textual descriptions, enabling cross-domain generalization without requiring ground-truth depth at inference. However, RSA[59] is highly sensitive to textual variation—semantically identical but differently textual descriptions can lead to significant performance degradation. VGLD incorporates high-level semantic features from visual inputs in conjunction with textual descriptions, enabling more robust metric depth scale recovery that is less susceptible to linguistic ambiguity and misalignment.
2.3 Language Modality for Metric Depth Estimation
Recent advances in vision-language models[26, 27, 38, 35, 20], driven by large-scale pretraining across diverse datasets, have led to strong cross-modal representations. These models now serve as effective baselines for various downstream tasks[13, 57, 32, 51, 65] and have been actively explored in monocular depth estimation. DepthCLIP[61] was the first to apply CLIP[38] to this task, reformulating depth regression as a distance classification problem using natural language descriptions such as "This object is giant, close…far…". By leveraging CLIP’s[38] semantic priors, DepthCLIP[61] enables zero-shot depth prediction without additional training by classifying relative distances across image regions. Subsequent works have extended this framework to improve adaptability and performance. Auty et al.[2] introduced learnable prompts to replace fixed textual tokens, enhancing the model’s flexibility and effectiveness. Hu et al.[18] further proposed a depth codebook with learnable prompts to better handle domain shifts in various scenes. CLIP2Depth[22] introduced mirror embeddings—non-natural language representations—to adapt CLIP[38] for depth estimation while avoiding the need for explicit text prompts. In parallel, methods such as VPD[63] , TADP[23] , EVP[24] and GeoWizard[11] extract high-level semantics from pretrained text-to-image diffusion models to enhance downstream tasks like depth estimation. Among these, Wordepth[58] was the first to model language as a variational prior for monocular depth estimation. It explicitly encodes object attributes—such as location, size, and orientation—from textual descriptions to align scale-invariant depth predictions with real-world metric scales. Recently, RSA[59] proposed an explicit constraint that enables metric depth recovery from relative estimates by predicting scale directly from textual descriptions. In contrast, VGLD leverages visual features from the CLIP image encoder in conjunction with textual descriptions, yielding more stable scale predictions and improved metric depth estimation over text-only baselines.
3 Method
3.1 Preliminaries
@@
The objective of monocular depth estimation is to predict continuous per-pixel depth values from a single RGB image[9]. We consider a dataset $\mathcal{D}=\{(I^{(n)},t^{(n)},d^{(n)}_{gt},dm^{(n)}_{gt})\}_{n=1}^{N}$ consisting of $N$ samples, where each sample includes an RGB image $I\in\mathbb{R}^{3\times H\times W}$, a corresponding linguistic description $t$, a ground-truth metric depth map $d_{gt}\in\mathbb{R}^{H\times W}$ and a ground-truth domain labels $dm_{gt}\in\{0,1\}$ which represent indoor or outdoor scene. We build upon a pretrained monocular relative depth estimation model $h_{\theta}$, which serves as the foundation for our metric depth scale recovery framework. Given an RGB image, the model predicts an inverse relative depth map $x\in\mathbb{R}^{H\times W}$, which lacks absolute scale information. To recover metric-scale depth from this scaleless prediction, we apply a global linear transformation informed by both the linguistic description and high-level visual semantics of the image. Specifically, similar to RSA[59], we predict a pair of scalars $(\hat{k},\hat{b})\in\mathbb{R}^{2}$ that represent the scale and shift parameters of the transformation. The final metric depth prediction is then computed as:
@@
3.2 VGLD
To model the relationship between the linear transformation parameters and the semantic content of both the image and its linguistic description, we leverage the CLIP model as a feature extractor. Benefiting from large-scale contrastive pretraining[38], CLIP provides a shared latent space that is well-suited for aligning object-centric visual and linguistic representations. Given an input sample $\{I,t\}$, we first extract visual and text embeddings using the CLIP image encoder and CLIP text encoder, respectively. The resulting embeddings are concatenated to form a fused representation, which is subsequently passed through a lightweight encoder network, GlobalNet—a three-layer MLP—to produce a compact 256-dimensional latent embedding used for downstream scale parameter regression.
Following ZoeDepth[5], we employ a lightweight MLP-based classifier, referred to as the Domain Routing Mechanism (DRM), to predict the domain of the input image based on its latent embedding. We consider two domains: indoor and outdoor. The predicted domain is then used to route the latent embedding to the corresponding domain-specific scalars prediction head.
3.3 Loss Function
As illustrated in Figure 2, the VGLD model freezes the weights of both the CLIP backbone and the relative depth estimator during training, and updates only the parameters of the GlobalNet and DRM modules. These modules are jointly optimized under a unified loss function. Since VGLD focuses on predicting a pair of global scalars rather than pixel-wise metric depth values, we do not adopt the Scale-Invariant Logarithmic Loss, which is more suitable for dense depth estimation tasks. Instead, following RSA[59], we adopt the L1 loss, which provides a more direct and interpretable supervision signal for scalars regression. The $\mathcal{L}_{metric}$ is formulated as:
@@ 3.3 Loss Function
where $\hat{d}_{pred}$ denotes the predicted metric depth, $(i,j)\in\Omega$ represents the image coordinates, $m(\cdot)\in\{0,1\}$ denotes the binary mask map and $M$ represents the number of pixels with valid ground truth values. We introduce a domain classification loss, denoted as $\mathcal{L}_{\text{dm}}$, implemented using the cross-entropy loss:
@@
where $\hat{dm}_{pred}\in\{0,1\}$ is the predicted domain label, and $dm_{gt}\in\{0,1\}$ is the corresponding ground-truth domain. To guide the model towards the optimal solution, we employ an MSE loss to provide LM loss (scalars supervision) for the modules:
@@
where $(\hat{k},\hat{b})$ are the predicted LM scalars from VGLD, and $(k_{\text{lm}},b_{\text{lm}})$ are the corresponding pseudo-labels provided by the Levenberg-Marquardt algorithm. The total loss is defined as follows:

### v3

@@ Abstract
Monocular depth estimation can be broadly categorized into two directions: relative depth estimation, which predicts normalized or inverse depth without absolute scale, and metric depth estimation, which aims to recover depth with real-world scale. While relative methods are flexible and data-efficient, their lack of metric scale limits their utility in downstream tasks. A promising solution is to infer absolute scale from textual descriptions. However, such language-based recovery is highly sensitive to natural language ambiguity, as the same image may be described differently across perspectives and styles. To address this, we introduce VGLD (Visually-Guided Linguistic Disambiguation), a framework that incorporates high-level visual semantics to resolve ambiguity in textual inputs. By jointly encoding both image and text, VGLD predicts a set of global linear transformation parameters that align relative depth maps with metric scale. This visually grounded disambiguation improves the stability and accuracy of scale estimation. We evaluate VGLD on representative models, including MiDaS and DepthAnything, using standard indoor (NYUv2) and outdoor (KITTI) benchmarks. Results show that VGLD significantly mitigates scale estimation bias caused by inconsistent or ambiguous language, achieving robust and accurate metric predictions. Moreover, when trained on multiple datasets, VGLD functions as a universal and lightweight alignment module, maintaining strong performance even in zero-shot settings. Code will be released upon acceptance.
@@ Introduction
Monocular depth estimation is a fundamental and long-standing task in computer vision, with applications ranging from autonomous driving(Schön et al. 2021), augmented reality(Ganj et al. 2023) to 3D reconstruction(Mescheder et al. 2019). The goal is to predict dense depth maps from single RGB images. However, reconstructing 3D geometry from a single image is an ill-posed problem because perspective projection causes a loss of depth dimension: any point along a projection ray corresponds to the same image coordinate. Consequently, the absolute distance from the camera to the scene cannot be directly recovered from a single view. Without camera calibration, additional sensors (e.g., IMU(Wofk et al. 2023), LiDAR(Lin et al. 2024)), or strong priors such as pre-trained depth models, scale ambiguity arises. While stereo or multi-view images can resolve scale by localizing points in 3D space, modern monocular depth estimation models are often trained on diverse datasets with varying data types and distributions—including single RGB images, video streams, and images with or without calibration parameters. These differences exacerbate the challenge of scale ambiguity, especially when deploying models across domains such as indoor and outdoor scenes.
To address scale ambiguity in monocular depth estimation, one line of work trains on multi-domain datasets (e.g., indoor and outdoor) to learn depth from domain-specific distributions(Ranftl et al. 2020; Reiner et al. 2023; Yang et al. 2024a, b). However, dataset biases limit generalization(Piccinelli et al. 2024). An alternative strategy is to leverage complementary cues shared across domains. Recent approaches explore language as a modality to resolve scale ambiguity without requiring expensive sensors (e.g., LiDAR). RSA(Zeng et al. 2024b) pioneers this direction by hypothesizing that textual descriptions can guide scale estimation and demonstrates that scale-less relative depth can be mapped to metric predictions via a language-guided global transformation.
Nevertheless, linguistic inputs are inherently ambiguous—semantically similar captions may produce inconsistent scales (see Figure 1), affecting stability. Still, language is robust to visual challenges like lighting or occlusion.
To reduce linguistic ambiguity, we propose a Visually-Guided Linguistic Disambiguation (VGLD) framework, which enriches textual inputs with semantic features extracted from the corresponding image using a CLIP Image Encoder(Radford et al. 2021). Additionally, to handle cross-domain depth variation, we introduce a Domain Router Mechanism (DRM) inspired by ZoeDepth(Bhat et al. 2023), which routes inputs to domain-specific heads for consistent metric predictions. To further stabilize training, we formulate depth scale recovery as a scalar regression task and supervise it using pseudo-labels $(k_{lm},b_{lm})$ obtained via the Levenberg-Marquardt algorithm. This nonlinear optimization technique helps guide the model toward an accurate training trajectory, enhancing robust scale recovery.
Our contributions are as follows:
  • We integrate high-level semantic information from the corresponding image alongside the textual description, thereby stabilizing the output of the scalars parameters;
  • We introduce the Domain Router Mechanism, which aids in solving the cross-domain estimation problem;
  • We leverage the Levenberg-Marquardt algorithm to optimize the training trajectory and guide the model’s training process;
  • Extensive experiments demonstrate the effectiveness of our method in both indoor and outdoor scenarios, highlighting its robustness to textual variations and strong zero-shot generalization.
@@
Related Work
Monocular Depth Estimation
@@
Monocular Depth Estimation (MDE) is a fundamental task in computer vision, with its development generally following two main directions: relative depth estimation and metric depth estimation. The goal of metric depth estimation is to predict pixel-wise depth values in metric units (e.g., meters), and models are typically trained by minimizing the discrepancy between predicted and ground-truth depth maps. In contrast, relative depth estimation focuses on inferring the ordinal relationships between pixel pairs, without providing any information about scale or units. A notable early milestone in this field was Eigen et al.(Eigen, Puhrsch, and Fergus 2014), the first to apply Convolutional Neural Networks (CNNs) to MDE. More recent methods such as AdaBins (Bhat, Alhashim, and Wonka 2021), LocalBins(Bhat, Alhashim, and Wonka 2022) and Binsformer(Li et al. 2024) reformulate the depth regression problem as a classification task through depth discretization. Multi-task learning strategies have also been explored: GeoNet(Qi et al. 2018) integrates surface normal estimation, while AiT(Ning et al. 2023) incorporates instance segmentation, both to enhance depth prediction through joint training. MiDaS(Ranftl et al. 2020; Reiner et al. 2023) and Diversedepth(Yin et al. 2020) advances relative depth estimation by pretraining on a diverse mixture of datasets, achieving strong generalization across domains. In addition, diffusion-based(Viola et al. 2024; Zhang et al. 2024; Song et al. 2025) methods, such as DDP (Ji et al. 2023), Marigold (Ke et al. 2024), and GeoWizard (Fu et al. 2024), adapt powerful diffusion priors to the depth estimation task via fine-tuning, enabling significant performance gains.
Metric Depth Scale Recovery
Relative depth estimation models have emerged as strong backbones for many metric depth Scale Recovery tasks, owing to their impressive cross-domain generalization and robustness. Building on MiDaS(Ranftl et al. 2020), DPT(Ranftl, Bochkovskiy, and Koltun 2021) replaces the convolutional backbone with a Vision Transformer and adapts it to metric depth via fine-tuning on scale-annotated datasets. ZoeDepth(Bhat et al. 2023) further enhances this pipeline by introducing a powerful decoder with a metric bins module, enabling effective scale recovery through supervised fine-tuning. Depth Anything extends ZoeDepth(Bhat et al. 2023) by replacing the MiDaS(Ranftl et al. 2020) encoder with its own architecture, achieving implicit conversion from relative to metric depth.
Other methods like Metric3D(Hu et al. 2024a; Yin et al. 2023), zeroDepth(Guizilini et al. 2023) and UniDepth(Piccinelli et al. 2024) recover scale by leveraging or predicting camera intrinsics, while PromptDA(Lin et al. 2024) introduces a lightweight LiDAR prompt to guide metric estimation. RSA(Zeng et al. 2024b) proposes an alternative paradigm by aligning relative depth with metric scale using textual descriptions, enabling generalization without requiring ground-truth depth at inference. However, RSA(Zeng et al. 2024b) is sensitive to linguistic variations, where semantically similar but differently worded inputs may cause inconsistent predictions. In contrast, VGLD leverages visual semantics to guide linguistic disambiguation, enabling more robust and reliable scale recovery. By grounding ambiguous textual inputs in high-level visual context, it mitigates sensitivity to language variation and achieves consistent metric depth estimation across domains.
Language Modality for Metric Depth Estimation
Recent advances in vision-language models(Li et al. 2022; Radford et al. 2021; Jia et al. 2022), driven by large-scale pretraining, have enabled strong cross-modal representations and inspired new approaches in monocular depth estimation. DepthCLIP(Zhang et al. 2022) first applied CLIP(Radford et al. 2021) to this task by reformulating depth regression as distance classification using natural language descriptions such as ”This object is giant, close…far…”, enabling zero-shot depth prediction via CLIP’s semantic priors. Subsequent works improved adaptability in various ways: Auty et al.(Auty and Mikolajczyk 2023) introduced learnable prompts to replace fixed text tokens; Hu et al.(Hu et al. 2024b) employed codebooks to address domain shifts; and CLIP2Depth(Kim and Lee 2024) proposed mirror embeddings to eliminate reliance on explicit textual input. Other approaches such as VPD(Zhao et al. 2023) , TADP(Kondapaneni et al. 2024) , EVP(Lavreniuk et al. 2023) and GeoWizard(Fu et al. 2024) extract semantic priors from pretrained text-to-image diffusion models to support depth prediction.
Recently, Wordepth(Zeng et al. 2024a) modeled language as a variational prior by explicitly encoding object attributes (e.g., size, position) to align relative predictions with metric depth. RSA(Zeng et al. 2024b) introduced a direct constraint to recover metric scale from text, but suffers from sensitivity to linguistic variation. In contrast, VGLD combines CLIP-based visual semantics with textual input, offering more stable and robust scale predictions compared to purely language-based methods.
Figure 3: Sensitivity to variations in linguistic descriptions on the NYUv2 dataset. We focus on the estimation results under three different textual inputs (text1-3). As shown in the depth maps, the RSA method exhibits noticeable sensitivity to textual variations, leading to inconsistent predictions—particularly in the regions highlighted by orange boxes. In contrast, our proposed VGLD produces more stable and consistent depth estimates across different descriptions. Warmer colors (red) indicate closer distances, while cooler colors (blue) indicate farther distances.
[…]
Abs Rel $\downarrow$ RMSE $\downarrow$ D1 $\uparrow$ Abs Rel $\downarrow$ RMSE $\downarrow$ D1 $\uparrow$
ZoeDepth(Bhat et al. 2023) robust depth estimation $\ddagger$ 0.077 0.277 0.953 0.054 2.281 0.971
ZeroDepth(Guizilini et al. 2023) 0.074 0.269 0.954 0.053 2.087 0.968
Metric3Dv2(Hu et al. 2024a) 0.047 0.183 0.989 0.044 1.985 0.985
MiDas-1(Reiner et al. 2023) Least Squares 0.121 0.388 0.866 0.333 6.901 0.408
Levenberg Marquardt 0.056 0.218 0.969 0.091 3.373 0.925
RSA-N/K(Zeng et al. 2024b) 0.171 0.569 0.731 0.163 4.082 0.798
RSA-NK(Zeng et al. 2024b) 0.168 0.561 0.737 0.160 4.232 0.782
VGLD-N/K-T (Ours) 0.158 0.529 0.758 0.133 3.755 0.854
VGLD-N/K-I (Ours) 0.121 0.423 0.860 0.120 3.668 0.868
VGLD-N/K-TCI (Ours) 0.119 0.414 0.867 0.120 3.598 0.871
VGLD-NK-T (Ours) 0.159 0.526 0.751 0.130 3.744 0.844
VGLD-NK-I (Ours) 0.123 0.426 0.855 0.122 3.574 0.868
VGLD-NK-TCI (Ours) 0.120 0.414 0.863 0.120 3.559 0.874
MiDas-2(Ranftl et al. 2020) Least Squares 0.130 0.421 0.845 0.336 6.925 0.421
Levenberg Marquardt 0.094 0.330 0.916 0.155 4.190 0.809
VGLD-N/K-T (Ours) 0.180 0.596 0.688 0.194 5.030 0.709
VGLD-N/K-I (Ours) 0.154 0.524 0.775 0.183 4.842 0.942
VGLD-N/K-TCI (Ours) 0.151 0.507 0.789 0.178 4.806 0.748
VGLD-NK-T (Ours) 0.182 0.615 0.682 0.191 4.994 0.723
VGLD-NK-I (Ours) 0.155 0.520 0.776 0.184 4.808 0.740
VGLD-NK-TCI (Ours) 0.151 0.513 0.780 0.180 4.804 0.737
DAV2-vits(Yang et al. 2024b) Least Squares 0.122 0.392 0.866 0.330 6.737 0.423
Levenberg Marquardt 0.052 0.209 0.969 0.103 3.277 0.919
VGLD-N/K-T (Ours) 0.163 0.546 0.713 0.166 4.189 0.756
VGLD-N/K-I (Ours) 0.128 0.433 0.830 0.154 4.219 0.756
VGLD-N/K-TCI (Ours) 0.125 0.423 0.842 0.152 3.872 0.779
VGLD-NK-T (Ours) 0.161 0.539 0.714 0.164 4.287 0.752
VGLD-NK-I (Ours) 0.127 0.436 0.835 0.160 4.031 0.761
VGLD-NK-TCI (Ours) 0.127 0.434 0.835 0.153 3.980 0.772
DAV1-vits(Yang et al. 2024a) Least Squares 0.121 0.397 0.863 0.331 6.772 0.423
Levenberg Marquardt 0.057 0.230 0.967 0.112 3.375 0.897
RSA-N/K(Zeng et al. 2024b) 0.147 0.484 0.775 0.160 4.437 0.780
RSA-NK(Zeng et al. 2024b) 0.148 0.498 0.776 0.158 4.457 0.786
VGLD-N/K-T (Ours) 0.145 0.496 0.792 0.151 4.354 0.773
VGLD-N/K-I (Ours) 0.115 0.405 0.872 0.144 4.074 0.790
VGLD-N/K-TCI (Ours) 0.112 0.390 0.887 0.140 4.081 0.807
VGLD-NK-T (Ours) 0.142 0.483 0.787 0.148 4.293 0.781
VGLD-NK-I (Ours) 0.114 0.404 0.880 0.142 4.151 0.814
VGLD-NK-TCI (Ours) 0.112 0.392 0.883 0.136 4.008 0.816
Table 1: Quantitative Depth Comparison on the NYUV2 and KITTI Dataset. $\dagger$ In the Model column, MiDas-1 denotes Midas-V3.1-dpt_swin2_large_384, MiDas-2 denotes Midas-V3.0-dpt_large_384, DAV2-vits denotes Depth-Anything-V2-Small, and DAV1-vits denotes Depth-Anything-V1-Small. $\ddagger$ denotes the results of certain state-of-the-art (SOTA) absolute scale estimation models. $\ast$ In the Method column, “N” and “K” indicate models trained on the NYUv2 and KITTI datasets, respectively. For example, VGLD-N/K-TCI refers to VGLD-N-TCI when evaluated on NYUv2, and VGLD-K-TCI when evaluated on KITTI. Best results are in bold, second best are underlined.
Method
Preliminaries
@@
The objective of monocular depth estimation is to predict continuous per-pixel depth values from a single RGB image(Eigen, Puhrsch, and Fergus 2014). We consider a dataset $\mathcal{D}=\{(I^{(n)},t^{(n)},d^{(n)}_{gt},dm^{(n)}_{gt})\}_{n=1}^{N}$ consisting of $N$ samples, where each sample includes an RGB image $I\in\mathbb{R}^{3\times H\times W}$, a corresponding linguistic description $t$, a ground-truth metric depth map $d_{gt}\in\mathbb{R}^{H\times W}$ and a ground-truth domain labels $dm_{gt}\in\{0,1\}$ which represent indoor or outdoor scene. We build upon a pretrained monocular relative depth estimation model $h_{\theta}$, which serves as the foundation for our metric depth scale recovery framework. Given an RGB image, the model predicts an inverse relative depth map $x\in\mathbb{R}^{H\times W}$, which lacks absolute scale information. To recover metric-scale depth from this scaleless prediction, we apply a global linear transformation informed by both the linguistic description and high-level visual semantics of the image. Specifically, similar to RSA(Zeng et al. 2024b), we predict a pair of scalars $(\hat{k},\hat{b})\in\mathbb{R}^{2}$ that represent the scale and shift parameters of the transformation. The final metric depth prediction is then computed as:
@@
VGLD
To model the relationship between the linear transformation parameters and the semantic content of both the image and its linguistic description, we leverage the CLIP model as a feature extractor. Benefiting from large-scale contrastive pretraining(Radford et al. 2021), CLIP provides a shared latent space that is well-suited for aligning object-centric visual and linguistic representations. Given an input sample $\{I,t\}$, we first extract visual and text embeddings using the CLIP image encoder and CLIP text encoder, respectively. The resulting embeddings are concatenated to form a fused representation, which is subsequently passed through a lightweight encoder network, GlobalNet—a three-layer MLP—to produce a compact 256-dimensional latent embedding used for downstream scale parameter regression.
Following ZoeDepth(Bhat et al. 2023), we employ a lightweight MLP-based classifier, referred to as the Domain Routing Mechanism (DRM), to predict the domain of the input image based on its latent embedding. We consider two domains: indoor and outdoor. The predicted domain is then used to route the latent embedding to the corresponding domain-specific scalars prediction head.
Loss Function
As illustrated in Figure 2, the VGLD model freezes the weights of both the CLIP backbone and the relative depth estimator during training, and updates only the parameters of the GlobalNet and DRM modules. These modules are jointly optimized under a unified loss function. Since VGLD focuses on predicting a pair of global scalars rather than pixel-wise metric depth values, we do not adopt the Scale-Invariant Logarithmic Loss, which is more suitable for dense depth estimation tasks. Instead, following RSA(Zeng et al. 2024b), we adopt the L1 loss, which provides a more direct and interpretable supervision signal for scalars regression. The $\mathcal{L}_{metric}$ is formulated as:
@@
where $\hat{d}_{pred}$ denotes the predicted metric depth, $(i,j)\in\Omega$ represents the image coordinates, $m(\cdot)\in\{0,1\}$ denotes the binary mask map and $M$ represents the number of pixels with valid ground truth values.
To ensure correct routing to the domain-specific scalars prediction head, We introduce a domain classification loss, denoted as $\mathcal{L}_{\text{dm}}$, implemented using the cross-entropy loss:
@@
where $\hat{dm}_{pred}\in\{0,1\}$ is the predicted domain label, and $dm_{gt}\in\{0,1\}$ is the corresponding ground-truth domain.
To guide the model towards the optimal solution, we employ an MSE loss to provide LM loss (scalars supervision) for the modules:
@@
where $(\hat{k},\hat{b})$ are the predicted LM scalars from VGLD, and $(k_{\text{lm}},b_{\text{lm}})$ are the corresponding pseudo-labels provided by the Levenberg-Marquardt algorithm. We assign a higher weight (10x) to the scale term $k_{\text{lm}}$ because empirical observations show that the model is more sensitive to errors in scale prediction than in shift. This design choice helps stabilize training and ensures more accurate depth scaling.
The total loss is defined as follows: